OpenStudy (anonymous):

Find the common difference of the sequence shown. 1/6,1/3,1/2,2/3 , ... 1/12 1/6 1/3

OpenStudy (tkhunny):

1/3 - 1/6 = ?? 1/2 - 1/3 = ?? 2/3 - 1/2 = ??

OpenStudy (anonymous):

thank you